Three cards are drawn in succession from an ordinary deck of playing cards. Find the probability that the first card is a red ace, the second card is a ten or jack, and the third card is greater than 3 but less than 7 if the cards are drawn
a) without replacement
b) with replacement

Found 2 solutions by sudhanshu_kmr, seriousgr8:
Answer by sudhanshu_kmr(1152)


a)in case of without replacement
no. of red ace = 2
no. of ten and jack = 2*4 = 8
no. of card which is greater than 3 but less than 7 = 3*4 = 12
probability = 2/52 * 8/51 * 12/50

b) in case of with replacement
probability = 2/52 * 8/52 * 12/52



Answer by seriousgr8(1)

this is absolutely correct!!!
a)in case of without replacement
no. of red ace = 2
no. of ten and jack = 2*4 = 8
no. of card which is greater than 3 but less than 7 = 3*4 = 12
probability = 2/52 * 8/51 * 12/50
b) in case of with replacement
probability = 2/52 * 8/52 * 12/52